OpenStudy (anonymous):

The perimeter of a rectangular field is 288 yards. If the width of the field is 66 yards, what is its length?

OpenStudy (anonymous):

@Firejay5

OpenStudy (firejay5):

288= 2l+ 2*66 288 = 2l + 132 156 = 2*l L = 78yds